OpenStudy (anonymous):

If you have a 1% by mass solution of NaCl how do you find molality? and how do you estimate the van't hoff factor?

OpenStudy (anonymous):

@Photon336 @agent0smith @Directrix any of you know how to do this?

OpenStudy (photon336):

Van't hoff factor is how many moles of ions are produced. \[NaOH \rightarrow Na^{+} + OH^{-}\] One mole of Na and one mole of OH produced therefore van hoff factor i = 2.

OpenStudy (anonymous):

Yes but our instructor wants us to estimate van't hoff not just assume it's 2...there is a certain way to do it however i don't know that way

OpenStudy (photon336):

BTW i didn't assume anything. one step is to figure out whether you have a non electrolyte or an electrolyte. NaOH is an ionic compound, so it will completely dissociate in water into one mole of Na and one mole of OH

OpenStudy (anonymous):

I understand that you didn't assume it...but the question is to estimate the actual van't hoff...There is a certain way to find the actual van't hoff factor but i don't remember how to do it. and i can't find it anywhere

OpenStudy (photon336):

so if a fraction alpha of the solute dissolves into n ions. \[i = \alpha (n-1) +1 \]

OpenStudy (photon336):

we know that NaOH dissolves completely because it is an ionic compound and it produces 2 ions = n

OpenStudy (photon336):

sorry, the correct word was dissociates. the fraction is going to be 1 because it completely dissociates. \[i = 1(2-1)+1 = 1+1 = 2 \]